OpenStudy (anonymous):

PLEASE HELP?!?! :)) Find the x and y values at the minimum of 2x^2+xy+2y^2+3y+2 And find the x and y values at the maximum of -3x^2+2xy-4y^2+2y+7

OpenStudy (anonymous):

What does min and max mean in this context?

OpenStudy (tkhunny):

You will need partial derivative for both x and y. I'll do y. x+4y+3 You do x.

OpenStudy (anonymous):

4x+y

OpenStudy (ranga):

Find partial derivative with respect to x and then y and set it to 0. Solve for x and y. partial derivative for x is: 4x+y = 0 partial derivative for y is: x+4y+3 = 0 solve for x and y.

OpenStudy (anonymous):

for the first equation I got x=0 and y=0, and for the second one I got x=-3 and y=0

OpenStudy (ranga):

you solve the equations simultaneously.

OpenStudy (anonymous):

x=0 and y=-3/5

OpenStudy (ranga):

I am getting (1/5, -4/5)

OpenStudy (anonymous):

i see, can we go through the second one now

OpenStudy (ranga):

can you verify the answer to the first one?

OpenStudy (ranga):

can you verify the answer to the first one? does the book give you the answer?

OpenStudy (anonymous):

i believe it is correct, it is

OpenStudy (anonymous):

for the second part I got -6x+2y and 2x-8y+2

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

now you set those equations equal to each other?

OpenStudy (ranga):

each expression set to 0.

OpenStudy (anonymous):

i got y=3/11 and x=36/11

OpenStudy (ranga):

can you show the work?

OpenStudy (anonymous):

nvm i got x its .09091

OpenStudy (ranga):

I am getting x = 1/11 and y = 3/11

OpenStudy (ranga):

Strictly speaking you are supposed to take the second partial derivative and prove these points are minimum/maximum. But that is up to you.

OpenStudy (tkhunny):

I am getting (1/5, -4/5) You do NOT have a minimum or maximum established. You MUST find the three second derivatives and make sense of it.

OpenStudy (tkhunny):

I am getting x = 1/11 and y = 3/11 You do NOT have a minimum or maximum established. You MUST find the three second derivatives and make sense of it. I did all the work on the 1sty Derivatives. It's your turn. Please demonstrate the THREE second derivatives.